# Improving Composed Image Retrieval via Contrastive Learning with Scaling Positives and Negatives

### v1

@@
Abstract
The Composed Image Retrieval (CIR) task aims to retrieve target images using a composed query consisting of a reference image and a modified text. Advanced methods often utilize contrastive learning as the optimization objective, which benefits from adequate positive and negative examples. However, the triplet for CIR incurs high manual annotation costs, resulting in limited positive examples. Furthermore, existing methods commonly use in-batch negative sampling, which reduces the negative number available for the model. To address the problem of lack of positives, we propose a data generation method by leveraging a multi-modal large language model to construct triplets for CIR. To introduce more negatives during fine-tuning, we design a two-stage fine-tuning framework for CIR, whose second stage introduces plenty of static representations of negatives to optimize the representation space rapidly. The above two improvements can be effectively stacked and designed to be plug-and-play, easily applied to existing CIR models without changing their original architectures. Extensive experiments and ablation analysis demonstrate that our method effectively scales positives and negatives and achieves state-of-the-art results on both FashionIQ and CIRR datasets. In addition, our methods also perform well in zero-shot composed image retrieval, providing a new CIR solution for the low-resources scenario.
1 Introduction
@@
Figure 1: Task introduction and the motivation of this work.
@@
Composed Image Retrieval (CIR) aims to retrieve images given a query composed of a modified text and a reference image. Unlike the standard text-to-image retrieval tasks, the modified text in CIR describes the unsatisfied attributes of the reference image or the new attributes based on the reference image. CIR provides a new idea for iteratively optimizing the retrieval results based on the current text-to-image retrieval and thus has become a popular research task in the multi-modal field. Previous research on CIR typically involves model architecture Vo et al. (2019); Delmas et al. (2022); Baldrati et al. (2024) and optimization objectives Baldrati et al. (2024); Wen et al. (2023); Liu et al. (2023c); Bai et al. (2023). The methods for the model architecture focus on better representation and fusion methods for texts and images. The contribution of the works in this aspect includes (1) introducing vision-language pre-trained models, like CLIP Radford et al. (2021), BLIP Li et al. (2022), as the backbone Baldrati et al. (2024); Wen et al. (2023); Liu et al. (2023c); Bai et al. (2023) and (2) designing the novel late fusion Vo et al. (2019); Wen et al. (2023); Baldrati et al. (2024) or early fusion Levy et al. (2024); Liu et al. (2023c); Bai et al. (2023) modules to fuse the reference image and the modified text to obtain the single query representation. Therefore, the popular model architecture in CIR can be illustrated in Fig.1(a), which consists of a query encoder and a target image encoder. In practice, a collection of image candidates is first converted into image representations by the image encoder for rapid indexing. When the user gives a reference image and a modified text, they are forwarded to the query encoder to fusion and obtain the query representation. Finally, the query representation is computed with the dot product or cosine similarity with all representations of candidate images, and the image with the top similarity is considered the target image.
@@
The works for optimization objective Wen et al. (2023); Bai et al. (2023) focus on aligning the query representation with the target image representations. Advanced methods use contrastive learning van den Oord et al. (2018) to align the representations of queries and images. The key to contrastive learning is to select correct and sufficient positives and negatives. The annotated triplets in the dataset, in the form of (reference image, modified text, target image), usually are regarded as positive examples, while negative examples are generated by replacing the target images with other ones in the mini-batch. However, as shown in Fig.1(b), there are two challenges with these works: (1) The number of manually annotated triplets (2̃0K) is deficient, leading to a lack of sufficient positive examples for the model. As a comparison, in other tasks using contrastive learning like visual representation learning Deng et al. (2009), image-text retrieval Lin et al. (2014); Plummer et al. (2015), and image retrieval Ong et al. (2017), the number of positive examples is at least 60k; (2) Previous CIR tasks typically use in-batch negative sampling, with around 128 negative examples, while many successful works in contrastive learning use over 4k negative examples Chen et al. (2020); He et al. (2020); Li et al. (2022). Existing works ignore these two problems at the data level, resulting in the inability of contrastive learning to fulfill its capabilities.
@@
Therefore, this work is based on a universal and simple motivation: to scale the number of positive and negative samples of the CIR task to the same scale as other tasks with contrastive learning. To construct more positives for CIR, we propose a novel data generation method based on the multi-modal Large Language Model (MLLM). Specifically, we design a four-step pipeline to automatically construct positive samples, which includes (1) caption generation with MLLM; (2) reference-target image pair matching; (3) modified text generation based on templates; and (4) positive example construction. With the help of our method, plenty of acceptable positive examples can be generated without any manual annotation, scaling the triplet number from 20k to 100k without the use of external datasets (Fig.1(b)). To introduce more negatives for CIR, we design a two-stage fine-tuning framework. Specifically, in the first stage, we follow previous works Baldrati et al. (2022); Wen et al. (2023); Liu et al. (2023c); Bai et al. (2023) and use in-batch negative sampling to enable the model to learn initial representation space for CIR; while in the second stage, we initialize the model trained in the first stage and freeze the target image encoder, only fine-tuning the query encoder. The frozen target image encoder introduces a large number of static representations of negatives at once (Fig.1(b)), guiding query encoders to optimize representation space rapidly. Note that the second stage has only about 1/20 of the time overhead of the first stage and can be easily superimposed on existing advanced models in CIR.
@@
  • We propose a two-stage plug-to-play framework to scale negative examples during fine-tuning, whose second stage can be quickly adapted to almost any model in CIR with 1/20 time overhead of the first stage.
@@ 2 Related Work
The recent paradigm in CIR Vo et al. (2019); Zhang et al. (2022); Zhao et al. (2022); Baldrati et al. (2022) consists of three main steps: (1) extracting the representation of both images and sentences; (2) fusing the representations of sentences and reference images to obtain query representations; (3) aligning the representations of queries and target images with similar semantics. For the first step, early models in CIR utilize two separate encoders Vo et al. (2019); Dodds et al. (2020); Kim et al. (2021); Xu et al. (2023a) while recent CIR models Liu et al. (2021); Zhao et al. (2022); Baldrati et al. (2022); Wen et al. (2023) exploit pre-trained vision-language encoders Lu et al. (2019); Radford et al. (2021) as the backbone. Some works Zhao et al. (2022); Baldrati et al. (2022) simply use the global representations extracted from these pre-trained encoders while other works Xu et al. (2023a); Wen et al. (2023); Zhang et al. (2022) integrate local and global representations. For the second step, some works Vo et al. (2019); Zhao et al. (2022); Baldrati et al. (2022); Wang et al. (2022) leverage weights or gating mechanisms, while other works Xu et al. (2023a); Zhang et al. (2022) design combining modules like cross-modal transformer. For the third step, the most commonly used loss functions in CIR are triplet loss Schroff et al. (2015); Zhang et al. (2022); Chen and Bazzani (2020), contrastive learning van den Oord et al. (2018); Vo et al. (2019); Baldrati et al. (2022); Lee et al. (2021); Zhao et al. (2022); Wen et al. (2023); Wang et al. (2022). Recent advanced methods in the CIR predominantly employ a combination of dual encoders and contrastive learning with in-batch negative sampling. We treat the models obtained from these methods as the first-stage models and continue to train them in the second stage to improve CIR performance further.
@@
InstructPix2Pix Brooks et al. (2023) first uses GPT-3 to generate modified text for captions and then utilizes a diffusion model to generate images for these texts. COVR Ventura et al. (2024) mine similar captioned videos from a large database and use a language model to generate modified text that describes the differences between the videos, resulting in the WebVid-CoVR dataset with 1.6 million triplets. CASE Levy et al. (2024) uses a data roaming approach that rephrases labels from a large-scale VQA dataset into a form suitable for composed image retrieval. CompDiff Gu et al. (2023) constructs triplets for CIR datasets by automatically generating modified texts and corresponding images using large language models and diffusion models. Unlike these works that often require generating images or well-labeled datasets, our method is built on a real image collection without the need for any additional manual annotation and leverages the capabilities of MLLM to construct triplets.
@@
In the realm of contrastive learning, negative sampling techniques have evolved to enhance model performance: the in-batch negative sampling from SimCLR Chen et al. (2020) selects negative examples from the same batch, while the memory bank approach in Bank Wu et al. (2018) utilizes a stored set of past instances for more diverse negatives. Additionally, MoCo He et al. (2020) employs a moving average of representations to create dynamic negatives, contributing to robust representation learning. Compared to Memory Bank and MoCo, our method does not dynamically update the negatives with the aid of additional queues or momentum encoders; instead, it fine-tunes the model for the second stage by introducing a large number of static negative samples at once.
3 Method
3.1 Preliminary
@@
$\mathcal{L}_{\rm cl}^{\rm t}=\frac{1}{B}\sum_{i=1}^{B}-\log(\frac{\exp(f(\mathbf{q}_{i},\mathbf{t}_{i})/\tau)}{\sum_{j=1}^{B}\exp(f(\mathbf{q}_{i},\mathbf{t}_{j})/\tau)})$ (2)
@@
3.2 Scaling Positive Examples
Due to the high cost of manually labeling triplets, we propose a simple but effective method with a multi-modal Large Language Model (MLLM) to construct the triplets for CIR. As shown in Fig,2, given an image dataset [^2] Image dataset here could be $\Omega$ in the CIR dataset or any image dataset. $D=\{I_{1},I_{2},...,I_{M}\}$ with size $M$, our method consists of four steps: (1) Generating a suitable caption for each image to obtain the image-text pairs; (2) Constructing M (reference image, target image) pairs; (3) Generating modified texts for image pairs using the captions; (4) Combining the modified texts and image pairs to form triplets.
Figure 2: Overview of Our Framework of Scaling Positive Examples and Negative Examples. We abbreviate some of the modified texts due to space constraints.
@@ 3.2 Scaling Positive Examples
Note that we attempt to use LLM to post-process the generated modified text. The first method involves using LLM to make the modified text more diverse and fluent. The second method uses in-context learning to make LLM mimic the modified text in annotated datasets. However, based on our experiments, neither of these methods surpasses the prompt template method. Specific results can be found in supplementary materials.
@@
Finally, we could combine image pairs from the second step with the modified texts obtained in the third step to get new M triplets $\{(I_{i},m_{i}^{\rm temp},I_{i}^{\rm t})\}$. So, we can obtain an expanded dataset that is comparable in size to the original dataset. We could use these new examples as a complement to the annotated dataset. We could also use these examples to train a model from scratch, thus allowing for fully automated training of a CIR model without human involvement.
3.3 Scaling Negative Examples
Recent works in visual contrastive representation learning Wu et al. (2018); He et al. (2020) have shown that scaling negative numbers can effectively improve performance. However, existing works in CIR employ in-batch negative sampling strategies, restricting the model from seeing enough negatives. Furthermore, recalling that the labeled data in CIR is a triplet, it is theoretically possible to construct negative examples by replacing any element in the triplet. Most works Vo et al. (2019); Baldrati et al. (2024); Wen et al. (2023); Bai et al. (2023) only use the "replace the target image" strategy to construct negative samples without additional interpretation. Therefore, we first explore the performance impact of different methods of constructing negative examples and find that "Replacing the target image" leads to more true and hard negatives than other methods with the popular CIR datasets. After determining the method of negative example construction, we propose a two-stage fine-tuning strategy for CIR that leverages a two-stage framework to scale negative examples during fine-tuning.
Figure 3: Performance of four different methods on negative example construction. The number on the horizontal axis corresponds to the serial number before the different replacing methods in Section 3.3.
@@ 3.3 Scaling Negative Examples
Most previous works use only the third method Vo et al. (2019); Delmas et al. (2022); Baldrati et al. (2024); Wen et al. (2023); Bai et al. (2023); Liu et al. (2023c), and Wang et al. (2022) uses the first three methods jointly. However, none of the existing works have explored all four methods completely. To this end, we compare these four negative construction methods while other settings remain the same. As shown in Fig.3, we find that constructing the negative examples by replacing target images works best. Based on the examples in Fig.2, we can observe that the other three methods easily generate relatively simple or false negatives. For example, since some modified texts (e.g., "a dog") only describe the target image, replacing the reference image with another image can lead to false negatives. Similarly, if the reference image is very similar to the target images, this type of data leads the model to directly use the reference image to retrieve the target image, making it easy to generate false negatives when replacing the modified text (e.g., "dog" and "lying"). Lastly, replacing the whole query pair leads to the simple negatives as the reference image and modified text significantly differ from those in the positive example (e.g., "sofa+shelf" and "llama+dog"). Compared with the other three methods, "replacing the target image" is inherently aligned with the final application scenario, and the probability of generating false negatives is relatively low. In the supplementary materials, we report the performance of every combination of four types of negative examples. The experimental results suggest that incorporating other types of negative examples may lead to increased overhead and potentially compromise model performance. For this reason, we keep consistent with previous work and only consider the negative example type of replacing the target image.
@@
Previous work on extended negative examples, such as Memory Bank Wu et al. (2018) and MoCo He et al. (2020), has focussed on visual representation learning, whose models typically follow a simple Siamese network architecture. However, CIR tasks require an information fusion of visual and language, and different methods follow different backbones, such as CLIP Radford et al. (2021), BLIP Li et al. (2022), and BLIP-2 Li et al. (2023). Therefore, we propose a more general two-stage framework to ensure fast adaptation of different models in CIR (Right half of Fig.2). Specifically, in the first stage, we fine-tune both the query encoder and the target encoder with in-batch negative sampling as in Eqn.2 following previous works Baldrati et al. (2024); Liu et al. (2023c); Bai et al. (2023); while in the second stage, we freeze the target encoder and only fine-tune the query encoder. Therefore, all candidate images, i.e., the entire $\Omega$, can advance past the frozen target image encoder, cached before the second fine-tuning stage. Finally, for triplet $(m_{i},r_{i},t_{i})$, we utilize all non-target images from the candidate set, i.e. $\Omega-\{t_{i}\}$, to form negative examples. The contrastive loss for the second stage can be expressed as
$\mathcal{L}^{\rm 2rd}_{\rm cl}=\frac{1}{B}\sum_{i=1}^{B}-\log\frac{\exp(f\left(\mathbf{q}_{i},~{}\hat{g}(t_{i})\right)/\tau)}{\sum_{t_{j}\in\Omega}\exp\left(f\left({\bf q_{i}},~{}\hat{g}(t_{j})\right)/\tau\right)}$ (7)
@@
4 Experiments
4.1 Experimental Setup
4.1.1 Baselines
To evaluate the superiority of our method, we conduct experiments on four advanced models in CIR: TG-CIR Wen et al. (2023), CLIP4CIR Baldrati et al. (2024), BLIP4CIR Liu et al. (2023c) and SPRC Bai et al. (2023).

### v2

@@
Abstract.
The Composed Image Retrieval (CIR) task aims to retrieve target images using a composed query consisting of a reference image and a modified text. Advanced methods often utilize contrastive learning as the optimization objective, which benefits from adequate positive and negative examples. However, the triplet for CIR incurs high manual annotation costs, resulting in limited positive examples. Furthermore, existing methods commonly use in-batch negative sampling, which reduces the negative number available for the model. To address the problem of lack of positives, we propose a data generation method by leveraging a multi-modal large language model to construct triplets for CIR. To introduce more negatives during fine-tuning, we design a two-stage fine-tuning framework for CIR, whose second stage introduces plenty of static representations of negatives to optimize the representation space rapidly. The above two improvements can be effectively stacked and designed to be plug-and-play, easily applied to existing CIR models without changing their original architectures. Extensive experiments and ablation analysis demonstrate that our method effectively scales positives and negatives and achieves state-of-the-art results on both FashionIQ and CIRR datasets. In addition, our method also performs well in zero-shot composed image retrieval, providing a new CIR solution for the low-resources scenario. Our code and data are released at https://github.com/BUAADreamer/SPN4CIR.
composed image retrieval, contrastive learning
copyright: acmlicensedjournalyear: 2024doi: 10.1145/3664647.3680808conference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.booktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28–November 1, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10ccs: Information systems Multimedia and multimodal retrievalccs: Information systems Image searchccs: Information systems Retrieval effectiveness
1. Introduction
@@
Figure 1. Task introduction and the motivation of this work.
[…]
[Introduction and Motivation]Task introduction and the motivation of this work.
@@
Composed Image Retrieval (CIR) aims to retrieve images given a query composed of a modified text and a reference image. Unlike the standard text-to-image retrieval tasks, the modified text in CIR describes the unsatisfied attributes of the reference image or the new attributes based on the reference image. CIR provides a new idea for iteratively optimizing the retrieval results based on the current text-to-image retrieval and thus has become a popular research task in the multi-modal field. Previous research on CIR typically involves model architecture (Vo et al., 2019; Delmas et al., 2022; Baldrati et al., 2024) and optimization objectives (Baldrati et al., 2024; Wen et al., 2023; Liu et al., 2023c; Bai et al., 2023). The methods for the model architecture focus on better representation and fusion methods for texts and images. The contribution of the works in this aspect includes (1) introducing vision-language pre-trained models, like CLIP (Radford et al., 2021), BLIP (Li et al., 2022), as the backbone (Baldrati et al., 2024; Wen et al., 2023; Liu et al., 2023c; Bai et al., 2023) and (2) designing the novel late fusion (Vo et al., 2019; Wen et al., 2023; Baldrati et al., 2024) or early fusion (Levy et al., 2024; Liu et al., 2023c; Bai et al., 2023) modules to fuse the reference image and the modified text to obtain the single query representation. Therefore, the popular model architecture in CIR can be illustrated in Fig.1(a), which consists of a query encoder and a target image encoder. In practice, a collection of image candidates is first converted into image representations by the image encoder for rapid indexing. When the user gives a reference image and a modified text, they are forwarded to the query encoder to fusion and obtain the query representation. Finally, the query representation is computed with the dot product or cosine similarity with all representations of candidate images, and the image with the top similarity is considered the target image.
@@
The works for optimization objective (Wen et al., 2023; Bai et al., 2023) focus on aligning the query representation with the target image representations. Advanced methods use contrastive learning (van den Oord et al., 2018) to align the representations of queries and images. The key to contrastive learning is to select correct and sufficient positives and negatives. The annotated triplets in the dataset, in the form of (reference image, modified text, target image), are usually regarded as positive examples, while negative examples are generated by replacing the target images with other ones in the mini-batch. However, as shown in Fig.1(b), there are two challenges with these works: (1) The number of manually annotated triplets (around 20K) is deficient, leading to a lack of sufficient positive examples for the model. As a comparison, in other tasks using contrastive learning like visual representation learning (Deng et al., 2009), image-text retrieval (Lin et al., 2014; Plummer et al., 2015), and image retrieval (Ong et al., 2017), the number of positive examples is at least 60k; (2) Previous CIR tasks typically use in-batch negative sampling, with around 128 negative examples, while many successful works in contrastive learning use over 4k negative examples (Chen et al., 2020; He et al., 2020; Li et al., 2022). Existing works ignore these two problems at the data level, resulting in the inability of contrastive learning to fulfill its capabilities.
@@
Therefore, this work is based on a universal and simple motivation: to scale the number of positive and negative samples of the CIR task to the same scale as other tasks with contrastive learning. To construct more positives for CIR, we propose a novel data generation method based on the multi-modal Large Language Model (MLLM). Specifically, we design a four-step pipeline to automatically construct positive samples, which includes (1) caption generation with MLLM; (2) reference-target image pair matching; (3) modified text generation based on templates; and (4) positive example construction. With the help of our method, plenty of acceptable positive examples can be generated without manual annotation, scaling the triplet number from 20k to 100k without the use of external datasets (Fig.1(b)). To introduce more negatives for CIR, we design a two-stage fine-tuning framework. Specifically, in the first stage, we follow previous works (Baldrati et al., 2022; Wen et al., 2023; Liu et al., 2023c; Bai et al., 2023) and use in-batch negative sampling to enable the model to learn the initial representation space for CIR; while in the second stage, we initialize the model trained in the first stage and freeze the target image encoder, only fine-tuning the query encoder. The frozen target image encoder introduces a large number of static representations of negatives at once (Fig.1(b)), guiding query encoders to rapidly optimize the representation space. Note that the second stage has only about 1/20 of the time overhead of the first stage and can be easily superimposed on existing advanced models in CIR.
@@
  • We propose a two-stage plug-and-play framework to scale negative examples during fine-tuning, whose second stage can be quickly adapted to almost any model in CIR with 1/20 time overhead of the first stage.
@@ 2. Related Work
The recent paradigm in CIR (Vo et al., 2019; Zhang et al., 2022; Zhao et al., 2022; Baldrati et al., 2022) consists of three main steps: (1) extracting the representation of both images and sentences; (2) fusing the representations of sentences and reference images to obtain query representations; (3) aligning the representations of queries and target images with similar semantics. For the first step, early models in CIR utilize two separate encoders (Vo et al., 2019; Dodds et al., 2020; Kim et al., 2021; Xu et al., 2023a) while recent CIR models (Liu et al., 2021; Zhao et al., 2022; Baldrati et al., 2022; Wen et al., 2023) exploit pre-trained vision-language encoders (Lu et al., 2019; Radford et al., 2021) as the backbone. Some works (Zhao et al., 2022; Baldrati et al., 2022) simply use the global representations extracted from these pre-trained encoders, while other works (Xu et al., 2023a; Wen et al., 2023; Zhang et al., 2022) integrate local and global representations. For the second step, some works (Vo et al., 2019; Zhao et al., 2022; Baldrati et al., 2022; Wang et al., 2022) leverage weights or gating mechanisms, while other works (Xu et al., 2023a; Zhang et al., 2022) design combining modules like cross-modal transformer. For the third step, the most commonly used loss functions in CIR are triplet loss (Schroff et al., 2015; Zhang et al., 2022; Chen and Bazzani, 2020), contrastive learning (van den Oord et al., 2018; Vo et al., 2019; Baldrati et al., 2022; Lee et al., 2021; Zhao et al., 2022; Wen et al., 2023; Wang et al., 2022). Recent advanced methods in the CIR predominantly employ a combination of dual encoders and contrastive learning with in-batch negative sampling. We treat the models obtained from these methods as the first-stage models and continue to train them in the second stage to further improve CIR performance.
@@
InstructPix2Pix (Brooks et al., 2023) first uses GPT-3 to generate modified text for captions and then utilizes a diffusion model to generate images for these texts. COVR (Ventura et al., 2024) mine similar captioned videos from a large database and use a language model to generate modified text that describes the differences between the videos, resulting in the WebVid-CoVR dataset with 1.6 million triplets. CASE (Levy et al., 2024) uses a data roaming approach that rephrases labels from a large-scale VQA dataset into a form suitable for composed image retrieval. CompDiff (Gu et al., 2023) constructs triplets for CIR datasets by automatically generating modified texts and corresponding images using large language models and diffusion models. Unlike these works that often require generating images or well-labeled datasets, our method is built on a real image collection without the need for any additional manual annotation and leverages the capabilities of MLLM to construct triplets.
@@
In the realm of contrastive learning, negative sampling techniques have evolved to enhance model performance: the in-batch negative sampling from SimCLR (Chen et al., 2020) selects negative examples from the same batch, while the memory bank approach in Bank (Wu et al., 2018) utilizes a stored set of past instances for more diverse negatives. Additionally, MoCo (He et al., 2020) employs a moving average of representations to create dynamic negatives, contributing to robust representation learning. Compared to Memory Bank and MoCo, our method does not dynamically update the negatives with the aid of additional queues or momentum encoders; instead, it fine-tunes the model for the second stage by introducing a large number of static negative samples at once.
3. Method
3.1. Preliminary
@@
(2) $\mathcal{L}_{\rm cl}^{\rm t}=\frac{1}{B}\sum_{i=1}^{B}-\log(\frac{\exp(f(\mathbf{q}_{i},\mathbf{t}_{i})/\tau)}{\sum_{j=1}^{B}\exp(f(\mathbf{q}_{i},\mathbf{t}_{j})/\tau)})$
@@
3.2. Scaling Positive Examples
Due to the high cost of manually labeling triplets, we propose a simple but effective method with a multi-modal Large Language Model (MLLM) to construct the triplets for CIR. As shown in Fig.2, given an image dataset [^2] Image dataset here could be $\Omega$ in the CIR dataset or any image dataset. $D=\{I_{1},I_{2},...,I_{M}\}$ with size $M$, our method consists of four steps: (1) Generating a suitable caption for each image to obtain the image-text pairs; (2) Constructing M (reference image, target image) pairs; (3) Generating modified texts for image pairs using the captions; (4) Combining the modified texts and image pairs to form triplets.
Figure 2. Overview of Our Framework of Scaling Positive Examples and Negative Examples. We abbreviate some of the modified texts due to space constraints.
[…]
[Overview]Overview of Our Framework of Scaling Positive and Negative Examples.
@@ 3.2. Scaling Positive Examples
Note that we also tried to use LLM to post-process the generated modified text. The first method involves using LLM to make the modified text more diverse and fluent. The second method uses in-context learning to make LLM mimic the modified text in annotated datasets. However, based on our experiments, neither of these methods surpasses the prompt template method. Specific results can be found in supplementary materials.
@@
Finally, we could combine image pairs from the second step with the modified texts obtained in the third step to get new M triplets $\{(I_{i},m_{i}^{\rm temp},I_{i}^{\rm t})\}$ [^3] Notably, since our image pair match step can match more number of target images for each image, our method can theoretically obtain up to $M^{2}$ triplets So, we can obtain an expanded dataset that is comparable in size to the original dataset. We could use these new examples as a complement to the annotated dataset. We could also use these examples to train a model from scratch, thus allowing for fully automated training of a CIR model without human involvement.
3.3. Scaling Negative Examples
Recent works in visual contrastive representation learning (Wu et al., 2018; He et al., 2020) have shown that scaling negative numbers can effectively improve performance. However, existing works in CIR employ in-batch negative sampling strategies, restricting the model from seeing enough negatives. Furthermore, recalling that the labeled data in CIR is a triplet, it is theoretically possible to construct negative examples by replacing any element in the triplet. Most works (Vo et al., 2019; Baldrati et al., 2024; Wen et al., 2023; Bai et al., 2023) only use the ”replacing the target image” strategy to construct negative samples without additional interpretation. Therefore, we first explore the performance impact of different methods of constructing negative examples and find that ”replacing the target image” leads to more true and hard negatives than other methods with the popular CIR datasets. After determining the method of negative example construction, we propose a novel fine-tuning strategy for CIR that leverages a two-stage framework to scale negative examples during fine-tuning.
Figure 3. Performance of four different methods on negative example construction. The number on the horizontal axis corresponds to the serial number before the different replacing methods in Section 3.3.
[…]
[Performance for different negative example types]Performance for different negative example types
@@ 3.3. Scaling Negative Examples
Most previous works use only the third method (Vo et al., 2019; Delmas et al., 2022; Baldrati et al., 2024; Wen et al., 2023; Bai et al., 2023; Liu et al., 2023c), and Wang et al. (2022) uses the first three methods jointly. However, none of the existing works have explored all four methods completely. To this end, we compare these four negative construction methods while other settings remain the same. As shown in Fig.3, we find that constructing the negative examples by replacing target images works best. Based on the examples in Fig.2, we can observe that the other three methods easily generate relatively simple or false negatives. For instance, since some modified texts (e.g., ”a dog”) only describe the target image, replacing the reference image with another image can lead to false negatives. Similarly, if the reference image is very similar to the target images, this type of data leads the model to directly use the reference image to retrieve the target image, making it easy to generate false negatives when replacing the modified text (e.g., ”dog” and ”lying”). Lastly, replacing the whole query pair leads to the simple negatives as the reference image and modified text significantly differ from those in the positive example (e.g., ”sofa+shelf” and ”llama+dog”). Compared with the other three methods, ”replacing the target image” is inherently aligned with the final application scenario, and the probability of generating false negatives is relatively low. In the supplementary materials, we report the performance of every combination of four types of negative examples. The experimental results suggest that incorporating other types of negative examples may lead to increased overhead and potentially compromise model performance. For this reason, we keep consistent with previous work and only consider the negative example type of ”replacing the target image”.
@@
Previous work on extended negative examples, such as Memory Bank (Wu et al., 2018) and MoCo (He et al., 2020), has focused on visual representation learning, whose models typically follow a simple Siamese network architecture. However, CIR tasks require an information fusion of visual and language, and different methods follow different backbones, such as CLIP (Radford et al., 2021), BLIP (Li et al., 2022), and BLIP-2 (Li et al., 2023). Therefore, we propose a more general two-stage framework to ensure fast adaptation of different models in CIR (Right half of Fig.2). Specifically, in the first stage, we fine-tune both the query encoder and the target encoder with in-batch negative sampling as in Eqn.2 following previous works (Baldrati et al., 2024; Liu et al., 2023c; Bai et al., 2023); while in the second stage, we freeze the target encoder and only fine-tune the query encoder. Therefore, all candidate images, i.e., the entire $\Omega$, can advance past the frozen target image encoder, cached before the second fine-tuning stage. Finally, for triplet $(m_{i},r_{i},t_{i})$, we utilize all non-target images from the candidate set, i.e. $\Omega-\{t_{i}\}$, to form negative examples. The contrastive loss for the second stage can be expressed as
(7) $\mathcal{L}^{\rm 2nd}_{\rm cl}=\frac{1}{B}\sum_{i=1}^{B}-\log\frac{\exp(f\left(\mathbf{q}_{i},~{}\hat{g}(t_{i})\right)/\tau)}{\sum_{t_{j}\in\Omega}\exp\left(f\left({\bf q_{i}},~{}\hat{g}(t_{j})\right)/\tau\right)}$
@@
4. Experiments
4.1. Experimental Setup
4.1.1. Baselines
To evaluate the superiority of our method, we conduct experiments on four advanced models in CIR: TG-CIR (Wen et al., 2023), CLIP4CIR (Baldrati et al., 2024), BLIP4CIR (Liu et al., 2023c) and SPRC (Bai et al., 2023).